An arithmetic sequence is given by 5, 8, 11, ....
Write down the value of d.
Find
I. u100
II. S100
Given that un = 1502, find the value of n.

Answer by stanbon(75887)

An arithmetic sequence is given by 5, 8, 11, ....
Write down the value of d.
----
d = 8-5 = 3
--------------------
Find
I. u(100) = u1+99d = 5+99*3
II. S(100) = (100/2)(5+5*99*3)
---------------------------------------
Given that un = 1502, find the value of n.
u(n) = 5 + (n-1)d = 1502
(n-1)3 = 1497
n-1 = 499
n = 500
-------------
Cheers,
Stan H.
------------